(5,1)is a solution to the equation 2x - 3y = 7. What does this mean?
Found 2 solutions by solver91311, mathsmiles:
Answer by solver91311(24713)



It means that if you substitute the coordinate of the given ordered pair for in the equation and the coordinate of the given ordered pair for in the equation, then the result will be a true statement. And indeed, 10 minus 3 is 7.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by mathsmiles(68)

When an answer is given in the form (5,1) and there is an x and y in the equation, the form is (x, y)

In this case, x=5 and y=1.

Checking:
2x - 3y = 7
2(5) - 3(1) = 7
10 - 3 = 7
7 = 7 Correct

Hope this help you with your other problems now too!!